# Low-Dimension-to-High-Dimension Generalization And Its Implications for Length Generalization

### v1

@@ 4 Main Results
We show theoretically how different models succeed or fail to achieve LDHD generalization as the effect of the inductive bias of the architectures under (S)GD. We focus on Boolean functions. More specifically, in the context of Boolean functions, we have $\Sigma=\{\pm 1\}$, ${\mathcal{X}}=\Sigma^{N}$, ${\mathcal{X}}_{n}=\Sigma^{n}\times\{1\}^{N-n}$ for $n=1,\dots,N$. The set of all Boolean functions potentially considered is ${\mathcal{F}}={\mathcal{F}}_{{\mathcal{X}},{\mathbb{R}}}$. We consider LDHD generalization from $N_{0}$ to $N$ for some $N_{0}<N$. Define $I(f)$ as the minimal set $I$ of indices that the function $f$ can be represented as a function of $x_{I}$, i.e., $I(f):=\operatorname*{arg\,min}_{I\subset[N]}|I|$ such that $f(x)=\tilde{f}(x_{I})$ for some function $\tilde{f}$ and all $x\in{\mathcal{X}}$. We say a function $f$ is $k$-sparse if $|I(f)|\leq k$.
Before presenting the theoretical results, we introduce two concepts degree profile w.r.t. independent set and min-degree interpolator w.r.t. independent set, which extend the concept degree profile and the concept min-degree interpolator, respectively. We use the two concepts to characterize the inductive bias of different model architectures under (S)GD.
Definition 2 (Degree Profile w.r.t. Independent Set ${\mathcal{B}}$).
Suppose that =${\mathcal{B}}=\{b_{1},\dots,b_{R}\}$ is an independent set of functions in ${\mathcal{F}}$ and $D=\max_{b\in{\mathcal{B}}}\deg(b)$. Let $f\in{\mathcal{F}}$ be a function in the subspace spanned by ${\mathcal{B}}$, i.e., $f=\sum_{i=1}^{R}\hat{f}_{{\mathcal{B}}}(b_{i})b_{i}$ for some $\hat{f}_{{\mathcal{B}}}(b_{i})\in{\mathbb{R}},i=1,\dots,R$. The degree profile of the function w.r.t. ${\mathcal{B}}$, denoted by $\operatorname{DegP}_{{\mathcal{B}}}(f)$, is a $(D+1)$-tuple where $D_{i}=\sum_{b\in{\mathcal{B}},\deg(b)=D+1-i}\hat{f}_{{\mathcal{B}}}(b)^{2}$ for $i=1,\dots,D+1$. The order of degree profiles is identical to the lexicographic order of the corresponding $D$-tuples.
Definition 3 (Min-Degree Interpolator w.r.t. Independent Set ${\mathcal{B}}$).
Suppose that =${\mathcal{B}}=\{b_{1},\dots,b_{R}\}$ is an independent set of functions in ${\mathcal{F}}$. Let ${\mathcal{X}}^{\prime}$ be a subset of the sample space ${\mathcal{X}}=\{\pm 1\}^{N}$. Denote the set of all interpolators on ${\mathcal{X}}^{\prime}$ for the concept $c$ by ${\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$, i.e., ${\mathcal{G}}_{{\mathcal{X}}^{\prime},c}=\{g\in{\mathcal{F}}\mid g(x)=c(x)\text{ for all }x\in{\mathcal{X}}^{\prime}\}$. A function $g$ is called the min-degree interpolator w.r.t. ${\mathcal{B}}$ on ${\mathcal{X}}_{0}$ for the concept $c$ if $g\in{\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$ and $\operatorname{DegP}(g)\leq\operatorname{DegP}(g^{\prime})$ for all $g^{\prime}\in{\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$.
@@ 4.1 Random Feature Model with Projection
We first consider the random feature model (RFM) and a class of its variants, i.e., Random Feature Models with Projection (RFMP; see Definition 4). The original RFM can be seen as a special instance of RFMP with $V=I_{N}$.
@@
where $K$ is the number of random features, $a=\left[a_{1},\dots,a_{K}\right]^{\intercal}$ is the learnable parameter, $w_{k}\sim{\mathcal{N}}(0,I_{r}/r)$, $b_{k}\sim{\mathcal{N}}(0,1/r)$ for $k=1,\dots,K$, and $\sigma$ is the activation function.
Technically, we follow the strongly expressive condition [1] for the activation function $\sigma$. Abbe et al. [1] shows that the RFM converges to the min-degree interpolator when initialized at 00 and trained with GD. However, this is not the case for all RFMP models. We show in Theorem 2 that an RFMP model converges to a min-degree interpolator w.r.t. an independent set determined by $V$.
@@
${\mathcal{B}}(V)=\left\{\chi^{V}_{T}(x)\right\}_{T\subseteq[r]},\quad\text{where }\chi^{V}_{T}(x)=\prod_{t\in T}\sum_{n=1}^{N}(v_{t})_{n}x_{n}$
Let $a_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $a$ is initialized at $a_{0}=0$ and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},c,V}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c^{*}(x)=c\left(\operatorname{Proj}(x,V)\right)$ that is $O_{N}(1)$-sparse. Then we have
$f_{\text{RFMP}}^{V,K}(x;a_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},c,V}}\operatorname{DegP}_{{\mathcal{B}}(V)}(g),\quad\text{as }K\to\infty,t\to\infty.$
When $V=I_{N}$, the independent set ${\mathcal{B}}(V)$ is the Fourier basis of the Boolean functions, and Theorem 2 implies that the RFM converges to the min-degree interpolator. From Theorem 2, we see that an RFMP model with the projection matrix $V$ can achieve LDHD generalization only if the target concept coincides with the min-degree interpolator w.r.t. ${\mathcal{B}}(V)$. Specially, for the RFM, we have:
@@
Corollary 1 shows that the min-degree interpolator and thus the RFM model can only achieve LDHD generalization for a very restricted set of functions that are only dependent on $x_{[N_{0}]}$. To achieve LDHD generalization with RFMP models requires prior knowledge of the concept class to design a projection to align the concepts with the min-degree interpolator w.r.t. the projected Fourier basis.
@@ 4.2 Position-Only Linear Attention with Advice
In this subsection, we investigate Position-Only Linear Attention with Advice (PLAA), which can be seen as a simplification of decoder-only Transformers (Definition 5), with a special focus position embeddings that are considered pivot to the length generalization of the Transformers [20, 11].
@@
where $A\in{\mathcal{U}}_{N}$ is the learnable parameter and $e_{n}$ denotes the vector with a $1$ in the $n$-th coordinate and 00’s elsewhere.
We further elaborate on the intuition behind the PLAA models. In the generation process of a decoder-only (linear) Transformer with position embeddings given input $\mathtt{s=s_{1}\dots s_{n}}$, the attention is computed by the query at the position $n$ and the keys at the positions $i\leq n$. The position of the query is special, advising the length of the input and reflecting the scale of the instance ideally. The PLAA model captures this feature and introduces the notation $n(x)$ to reflect the dimension of the subspace that $x$ belongs to. To further simplify and focus on the position embeddings, we assume that the value of each $x_{i}$ is identical to itself and the attention is only related to positions. The contribution of the interaction between the position embeddings is $\left[e_{1},\dots,e_{n}\right]^{\intercal}A_{[n],[n]}e_{n}=A_{[n],n}$ for some upper triangle matrix $A$. When the input is embedded to length $N$ but the query is still made at the position $n$, the output of the model is $x^{\intercal}Ae_{n(x)}$, i.e., the output of the PLAA model. Therefore, the PLAA model is a simplification of decoder-only Transformers that enables theoretical analysis of the impact of the position embeddings on length generalization.
In Definition 5, we directly parameterize the PLAA model directly with the attention matrix. In practice, however, the attention matrix is typically computed by the interaction between the position embeddings. Therefore, we consider the PLAA models with the Absolute Position Embedding (APE) and the Relative Position Embedding (RPE), respectively. See Definitions 6 and 7. Note that we consider Generalized RPE (GPRE) in Definition 7 because the RPE can be seen as a special instance of the GRPE with ${\mathcal{U}}={\mathcal{U}}_{\text{RPE}}=\{D_{1},\dots,D_{N}\}$, where $D_{k}$ is a $k$-th upper diagonal matrix such that $(D_{k})_{ij}$ is $1$ if $j=i+k-1$ and 00 otherwise. We seek a more general result applicable to all similar parameterization methods to the RPE.
@@ A PLAA model with APE is
where $M_{N}^{u}{\mathbb{R}}^{N\times N}$ is the upper triangle mask, i.e., $(M_{N}^{u})_{ij}$ is $1$ if $i\leq j$ and 00 otherwise, and $P\in{\mathbb{R}}^{d_{P}\times N}$ is the learnable parameter of the model.
@@
where $U_{i}\in{\mathcal{U}}_{N},i=1,\dots,r$ are upper triangle matrices that satisfy $\left\langle U_{i},U_{i}\right\rangle=1$ for all $i=1,\dots,r$ and $\left(U_{i}\right)_{kl}\left(U_{j}\right)_{kl}=0$ for all $i\neq j$ and $1\leq k,l\leq N$, and $p=\left[p_{1},\dots,p_{r}\right]^{\intercal}\in{\mathbb{R}}^{r}$ is the learnable parameter of the model.
@@
The condition $\left(U_{i}\right)_{kl}\left(U_{j}\right)_{kl}=0$ for all $i\neq j$ and $1\leq k,l\leq N$ in Definition 7 means each element in the position-only attention is characterized by at most one parameter. For example, in RPE, $A_{i,j}(i\leq j)$ is only parameterized by $p_{j-i}$. This condition naturally generalizes the intuition of RPE that characterizes some “relative” relations between two positions.
We first consider the vanilla PLAA model, the direct parameterization with the attention matrix. Theorem 3 shows that the PLAA model converges to the min-degree interpolator w.r.t. the independent set ${\mathcal{B}}_{N}^{\text{PLAA}}$ that is defined in the theorem.
@@
Let $A_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $A$ is initialized at $A_{0}=0$ and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},A^{*}}^{\text{PLAA}}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c(x)=f_{\text{PLAA}}(x;A^{*})$. Then we have
$f_{\text{PLAA}}(x;A_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},A^{*}}^{\text{PLAA}}}\operatorname{DegP}_{{\mathcal{B}}_{N}^{\text{PLAA}}}(g),\quad\text{as }t\to\infty.$
For the PLAA model with APE, Theorem 4 shows that the model converges to the interpolator whose attention matrix is of the minimal nuclear norm if the minimal-nuclear-norm-solution conjecture (Assumption 2 about the implicit regularization in matrix factorization holds. While the conjecture has not been theoretically justified in general cases, it is proved correct under some specific conditions and supported by empirical evidence [10, 4].
Assumption 2 (Minimal-Nuclear-Norm-Solution Conjecture [10]).
Consider the optimization
$\min_{U\in{\mathbb{R}}^{n\times n}}f(U)=\left\|{\mathcal{A}}\left(UU^{\intercal}\right)-y\right\|_{2}^{2},$
where ${\mathcal{A}}:{\mathbb{R}}^{n\times n}\mapsto{\mathbb{R}}^{m}$ is a linear operator specified by ${\mathcal{A}}(\cdot)_{i}=\left\langle A_{i},\cdot\right\rangle$ for symmetric linearly independent $A_{i}\in{\mathbb{R}}^{n\times n},1\leq i\leq m$, and $y\in{\mathbb{R}}^{m}$. Define $X=X(U):=UU^{\intercal}$. Let $U_{t}$ be the parameter reached by gradient flow in the timestep $t$ and $X_{t}$ be $U_{t}U_{t}^{\intercal}$. Define $X_{\infty}(X_{\text{init}}):=\lim_{t\to\infty}X_{t}$ for $U_{0}U_{0}^{\intercal}=X_{\text{init}}$. For any full rank $X_{0}$, if $\hat{X}=\lim_{\alpha\to 0}X_{\infty}(\alpha X_{0})$ exists and ${\mathcal{A}}\left(\hat{X}\right)=y$, then $\hat{X}\in\operatorname*{arg\,min}_{{\mathcal{X}}\succ\|X\|_{*}}$ s.t. ${\mathcal{A}}(X)=y$.
Theorem 4.
Suppose that Assumption 2 holds and $d_{P}=N$. Let $P_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $P$ is initialized at $P_{0}$ such that $(P_{0})_{ij}\sim{\mathcal{N}}(0,\alpha)$ for $(i,j)\in[d_{P}]\times[N]$ and $\alpha$ sufficiently small, and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},P^{*}}^{\text{APE}}$ be the set of all PLAA interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c(x)=f_{\text{PLAA}}^{\text{APE}}(x;P^{*})$. Then we have
$f_{\text{PLAA}}^{\text{APE}}(x;P_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},P^{*}}^{\text{APE}}}\|A_{g}\|_{*},\quad\text{as }t\to\infty,$
where $A_{g}$ is the attention matrix in the PLAA model $g$.
Theorem 4 shows that APE leads to a low-rank attention matrix in the PLAA model. Therefore, when the training data is restricted to ${\mathcal{X}}_{N_{0}}$, the rank of the attention matrix learned by the model with APE is no higher than $N_{0}$. This leads to length generalization failure when the target concept requires higher-rank attention to be represented. We have the following direct corollary of Theorem 4 that partially explains the limitation of APE for length generalization.
Corollary 2.
The PLAA with APE fails to achieve LDHD generalization for the concept $c(x)=f_{\text{PLAA}}(x;A^{*})$ if $\operatorname{rank}(A^{*})>N_{0}$.
Theorem 5 characterizes that the PLAA converges to the interpolator that minimizes the degree-profile w.r.t. to the independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$, which can be seen as the projections of ${\mathcal{U}}$ onto ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$. Theorem 3 is a special case of Theorem 5 for ${\mathcal{U}}=\left\{E_{ij}\right\}_{1\leq i\leq j\leq N}$, where $E_{ij}\in{\mathbb{R}}^{N\times N}$ is the elementary matrix that has a 1 at the $(i,j)$-th position and 0 everywhere else.
@@
Let $p_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $p$ is initialized at $p_{0}=0$ and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},p^{*}}^{\text{GRPE},{\mathcal{U}}}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c(x)=f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p^{*})$. Then we have
With the inductive bias of the PLAA with GRPE, Corollary 3 states that LDHD generalization can be achieved if and only if the target concept can be represented by the elements in ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ that have dependence on ${\mathcal{X}}_{N_{0}}$.
@@
Consider the concept $c(x)=f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p^{*})=\sum_{k=1}^{r}c_{k}\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x)$. Under the conditions of Theorem 5, the PLAA with GRPE achieves LDHD generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$ if and only if
@@
While the projection in the RFM and the position embeddings in the PLAA may introduce a stronger inductive bias that benefits LDHD generalization, they can also reduce the expressiveness of the models. The critical point is to align the models with the concept class in such a way that we can incorporate a strong inductive bias for LDHD generalization while maintaining sufficient expressiveness for the concept class.
5 Further Implications for Length Generalization
We discuss further implications of the LDHD generalization perspective for length generalization.

### v3

@@ 4 Main Results
We show theoretically how different models succeed or fail to achieve LDHD generalization as the effect of the inductive bias of the architectures under (S)GD. We focus on Boolean functions. More specifically, in the context of Boolean functions, we have $\Sigma=\{\pm 1\}$, ${\mathcal{X}}=\Sigma^{N}$, ${\mathcal{X}}_{n}=\Sigma^{n}\times\{1\}^{N-n}$ for $n=1,\dots,N$. The set of all Boolean functions potentially considered is ${\mathcal{F}}={\mathcal{F}}_{{\mathcal{X}},{\mathbb{R}}}$. Our analysis can be naturally extended to tasks over finite alphabets by considering their binary representations.
We consider LDHD generalization from $N_{0}$ to $N$ for some $N_{0}<N$. Define $I(f)$ as the minimal set $I$ of indices that the function $f$ can be represented as a function of $x_{I}$, i.e., $I(f):=\operatorname*{arg\,min}_{I\subset[N]}|I|$ such that $f(x)=\tilde{f}(x_{I})$ for some function $\tilde{f}$ and all $x\in{\mathcal{X}}$. A function $f$ is $k$-sparse if $|I(f)|\leq k$.
Before presenting the theoretical results, we introduce two concepts degree profile w.r.t. linearly independent set and min-degree interpolator w.r.t. linearly independent set, which extend the concept degree profile and the concept min-degree interpolator, respectively. We use the two concepts to characterize the inductive bias of different model architectures under (S)GD.
Definition 2 (Degree Profile w.r.t. Linearly Independent Set $\mathcal{B}$}).
Suppose that ${\mathcal{B}}=\{b_{1},\dots,b_{R}\}$ is a linearly independent set of functions in $\mathcal{F}$} and $D=\max_{b\in{\mathcal{B}}}\deg(b)$. Let $f\in{\mathcal{F}}$ be a function in the subspace spanned by $\mathcal{B}$}, i.e., $f=\sum_{i=1}^{R}\hat{f}_{{\mathcal{B}}}(b_{i})b_{i}$ for some $\hat{f}_{{\mathcal{B}}}(b_{i})\in{\mathbb{R}},i=1,\dots,R$. The degree profile of the function w.r.t. $\mathcal{B}$}, denoted by $\operatorname{DegP}_{{\mathcal{B}}}(f)$, is a $(D+1)$-tuple where $D_{i}=\sum_{b\in{\mathcal{B}},\deg(b)=D+1-i}\hat{f}_{{\mathcal{B}}}(b)^{2}$ for $i=1,\dots,D+1$. The order of degree profiles is identical to the lexicographic order of the corresponding $D$-tuples.
Definition 3 (Min-Degree Interpolator w.r.t. Linearly Independent Set $\mathcal{B}$}).
Suppose that ${\mathcal{B}}=\{b_{1},\dots,b_{R}\}$ is a linearly independent set of functions in $\mathcal{F}$}. Let $\mathcal{X}$$^{\prime}$ be a subset of the sample space ${\mathcal{X}}=\{\pm 1\}^{N}$. Denote the set of all interpolators on $\mathcal{X}$$^{\prime}$ for the concept $c$ by ${\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$, i.e., ${\mathcal{G}}_{{\mathcal{X}}^{\prime},c}=\{g\in{\mathcal{F}}\mid g(x)=c(x)\text{ for all }x\in{\mathcal{X}}^{\prime}\}$. A function $g$ is called the min-degree interpolator w.r.t. $\mathcal{B}$} on ${\mathcal{X}}_{0}$ for the concept $c$ if $g\in{\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$ and $\operatorname{DegP}(g)\leq\operatorname{DegP}(g^{\prime})$ for all $g^{\prime}\in{\mathcal{G}}_{{\mathcal{X}}^{\prime},c}$.
@@ 4.1 Random Feature Model with Projection
We first consider the random feature model (RFM) and a class of its variants, i.e., Random Feature Models with Projection (RFMP; see Definition 4). RFM is widely employed as approximations of practical neural network models in theoretical studies. By comparing the inductive biases introduced by RFM and RFMP under various projections, we demonstrate the importance of incorporating prior knowledge to achieve LDHD generalization and this prior knowledge can be effectively embedded through model design.
@@
where $K$ is the number of random features, $\sigma$ is the activation function, $a=\left[a_{1},\dots,a_{K}\right]^{\intercal}$ is the learnable parameter, and $w_{k}\sim{\mathcal{N}}(0,I_{r}/r)$, $b_{k}\sim{\mathcal{N}}(0,1/r)$ for all $k=1,\dots,K$.
The original RFM can be seen as a special instance of RFMP with $V=I_{N}$. Technically, we follow the strongly expressive condition (Abbe et al., 2023) for the activation function $\sigma$. Abbe et al. (2023) shows that the RFM converges to the min-degree interpolator when initialized at 0 and trained with GD. However, this is not the case for all RFMP models. We show in Theorem 2 that an RFMP model converges to a min-degree interpolator w.r.t. a linearly independent set determined by the set $V$.
@@
${\mathcal{B}}(V):=\left\{\chi^{V}_{T}(x)\right\}_{T\subseteq[r]},$
where
$\chi^{V}_{T}(x)=\prod_{t\in T}\sum_{n=1}^{N}(v_{t})_{n}x_{n}.$
Let $a_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $a$ is initialized at $a_{0}=0$ and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},c,V}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c^{*}(x)=c\left(\operatorname{Proj}(x,V)\right)$ that is $O_{N}(1)$-sparse. Then we have, as $K\to\infty,t\to\infty$,
$f_{\text{RFMP}}^{V,K}(x;a_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},c,V}}\operatorname{DegP}_{{\mathcal{B}}(V)}(g).$
When $V=I_{N}$, the linearly independent set ${\mathcal{B}}(V)$ is the Fourier basis of the Boolean functions, and Theorem 2 implies that the RFM converges to the min-degree interpolator. From Theorem 2, we see that an RFMP model with the projection matrix $V$ can achieve LDHD generalization only if the target concept coincides with the min-degree interpolator w.r.t. ${\mathcal{B}}(V)$. Specially, for the RFM, we have:
@@
Corollary 1 shows that the min-degree interpolator and thus the RFM model can only achieve LDHD generalization for a very restricted set of functions that are only dependent on $x_{[N_{0}]}$. Achieving LDHD generalization with RFMP models requires prior knowledge of the concept class to design the projection. Example 3 illustrates how LDHD generalization is possible for the target function with dependence beyond $x_{[N_{0}]}$ by choosing a proper projection.
Example 3.
Consider the target function $f(x)=4x_{1}+3x_{2}$, $N_{0}=1$, and $N=2$. The min-degree interpolator on ${\mathcal{X}}_{N_{0}}$ is $f_{1}(x)=4x_{1}$, which does not achieve LDHD generalization on ${\mathcal{X}}_{N}$. In the RFMP model, if we choose
$V=\begin{bmatrix}0.8&0.6\\ 0.6&-0.8\end{bmatrix}$,
then we have
$\begin{gathered}{\mathcal{B}}(V)=\{1,0.8x_{1}+0.6x_{2},0.6x_{1}-0.8x_{2},\\ (0.8x_{1}+0.6x_{2})(0.6x_{1}-0.8x_{2})\}.\end{gathered}$
The min-degree interpolator w.r.t. the linearly independent set ${\mathcal{B}}(V)$ on ${\mathcal{X}}_{N_{0}}$ is $f_{2}(x)=4x_{1}+3x_{2}=f(x)$, which achieves LDHD generalization on ${\mathcal{X}}_{N}$.
@@ 4.2 Position-Only Linear Attention with Advice
In this subsection, we investigate Position-Only Linear Attention with Advice (PLAA), which can be seen as a simplification of decoder-only Transformers (Definition 5), with a special focus position embeddings that are considered pivot to the length generalization of the Transformers (Shaw et al., 2018; Jelassi et al., 2023).
@@
where $A\in{\mathcal{U}}_{N}$ is the learnable parameter and $e_{n}$ denotes the vector with a $1$ in the $n$-th coordinate and 0’s elsewhere.
We further elaborate on the intuition behind the PLAA models. In the generation process of a decoder-only (linear) Transformer with position embeddings given input $\mathtt{s=s_{1}\dots s_{n}}$, the attention is computed by the query at the position $n$ and the keys at the positions $i\leq n$. The position of the query is special, advising the length of the input and reflecting the scale of the instance ideally. The PLAA model captures this feature and introduces the notation $n(x)$ to reflect the dimension of the subspace that $x$ belongs to. To further simplify and focus on the position embeddings, we assume that the value of each $x_{i}$ is identical to itself (i.e., we fix the value matrix in the attention to $I$ and thus $W_{V}x_{i}=x_{i}$) since the value head is not central to length generalization: the interpolator is expected to learn a suitable value head. We reasonably suppose a correct value head to focus on the specific challenge in length generalization. Additionally, we assume the attention is only related to positions, highlighting the standalone impact of position embeddings on inductive bias and length generalization. The contribution of the interaction between the position embeddings is $\left[e_{1},\dots,e_{n}\right]^{\intercal}A_{[n],[n]}e_{n}=A_{[n],n}$ for some upper triangle matrix $A$. When the input is embedded to length $N$ but the query is still made at the position $n$, the output of the model is $x^{\intercal}Ae_{n(x)}$, i.e., the output of the PLAA model. Therefore, the PLAA model is a simplification of decoder-only Transformers focusing on the impact of the position embeddings on length generalization. For a more detailed elaboration on PLAA, see Appendix B.
In Definition 5, we directly parameterize the PLAA model directly with the attention matrix. In practice, however, the attention matrix is typically computed by the interaction between the position embeddings. Therefore, we consider the PLAA models with the Absolute Position Embedding (APE) and the Relative Position Embedding (RPE), respectively. See Definitions 6 and 7. Note that we consider Generalized RPE (GPRE) in Definition 7 because the RPE can be seen as a special instance of the GRPE with ${\mathcal{U}}={\mathcal{U}}_{\text{RPE}}=\{D_{1},\dots,D_{N}\}$, where $D_{k}$ is a $k$-th upper diagonal matrix such that $(D_{k})_{ij}$ is $1$ if j=i+k1j=i+k-1 and 0 otherwise. We seek a more general result applicable to all similar parameterization methods to the RPE.
@@ A PLAA model with APE is
where $M_{N}^{u}\in{\mathbb{R}}^{N\times N}$ is the upper triangle mask, i.e., $(M_{N}^{u})_{ij}$ is $1$ if $i\leq j$ and 0 otherwise, and $P\in{\mathbb{R}}^{d_{P}\times N}$ is the learnable parameter of the model.
@@
where $U_{i}\in{\mathcal{U}}_{N},i=1,\dots,r$ are upper triangle matrices that satisfy $\left\langle U_{i},U_{i}\right\rangle=1$ for all $i=1,\dots,r$ and $\left(U_{i}\right)_{kl}\left(U_{j}\right)_{kl}=0$ for all $i\neq j$ and $1\leq k,l\leq N$, and $p=\left[p_{1},\dots,p_{r}\right]^{\intercal}\in{\mathbb{R}}^{r}$ is the learnable parameter.
@@
The condition $\left(U_{i}\right)_{kl}\left(U_{j}\right)_{kl}=0$ for all $i\neq j$ and $1\leq k,l\leq N$ in Definition 7 means each element in the position-only attention is characterized by at most one parameter. This condition generalizes the property of RPE that $A_{i,j}(i\leq j)$ is only parameterized by $p_{j-i}$.
For the PLAA model with APE, Theorem 3 shows that it converges to the min-degree interpolator w.r.t. the linearly independent set ${\mathcal{B}}_{N}^{\text{PLAA}}$.
@@
Suppose that $d_{P}\geq N$. Let $P_{t}(\alpha)$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $P$ is initialized at $P_{0}(\alpha)$ such that $P_{0}^{\intercal}(\alpha)P_{0}(\alpha)=\alpha I_{N}$ ($\alpha>0$), and optimized with gradient descent/gradient flow under $\ell_{2}$ loss (denoted by $L(P)$) on ${\mathcal{X}}_{N_{0}}$. Define $P_{\infty}(\alpha):=\lim_{t\to\infty}P_{t}(\alpha)$. Let ${\mathcal{G}}_{N_{0},A^{*}}^{\text{PLAA}}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c(x)=f_{\text{PLAA}}(x;A^{*})$. If $L\left((P_{\infty}(\alpha)\right)=0$ for all $0<\alpha\leq\alpha_{0}$ ($\alpha_{0}>0$ is some constant) and $\hat{P}:=\lim_{\alpha\to 0}P_{\infty}(\alpha)$, then
$f_{\text{PLAA}}^{\text{APE}}(x;\hat{P})=\arg\min_{g\in{\mathcal{G}}_{N_{0},A^{*}}^{\text{PLAA}}}\operatorname{DegP}_{{\mathcal{B}}_{N}^{\text{PLAA}}}(g).$
For any function $f$ that can be represented by a PLAA model, there exists a matrix $A^{f}\in{\mathbb{R}}^{N_{0}\times N_{0}}$ such that $f(x)=\sum_{1\leq i\leq j\leq N_{0}}A^{f}_{ij}b_{ij}^{\text{PLAA}}(x)$ for all $x\in{\mathcal{X}}_{N_{0}}$. Consequently, for $f^{\text{APE}}_{\text{PLAA}}(x)(x;\hat{P})=\sum_{1\leq i\leq j\leq N}\hat{A}_{ij}b_{ij}^{\text{PLAA}}(x)$, we have $\hat{A}_{ij}=0$ for all $j\geq i>N_{0}$. This implies that the PLAA with APE cannot achieve LDHD generalization for the concept $c(x)=f_{\text{PLAA}}(x;A^{*})$ if $A^{*}_{ij}\neq 0$ for some $j\geq i>N_{0}$, which partially explains the limitation of APE for length generalization.
The PLAA with GRPE can overcome the aforementioned limitation of the PLAA with APE in length generalization. Theorem 4 characterizes that the PLAA with GRPE converges to the interpolator that minimizes the degree-profile w.r.t. the linearly independent set ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$.
@@
Let $p_{t}$ be the learnable parameter at the timestep $t$ in the training process where the learnable parameter $p$ is initialized at $p_{0}=0$ and optimized with gradient descent/gradient flow under $\ell_{2}$ loss on ${\mathcal{X}}_{N_{0}}$. Let ${\mathcal{G}}_{N_{0},p^{*}}^{\text{GRPE},{\mathcal{U}}}$ be the set of all interpolators on ${\mathcal{X}}_{N_{0}}$ for the concept $c(x)=f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p^{*})$. Then we have, as $t\to\infty$,
$f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p_{t})\to\arg\min_{g\in{\mathcal{G}}_{N_{0},p^{*}}^{\text{GRPE},{\mathcal{U}}}}\operatorname{DegP}_{{\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}}(g).$
With the inductive bias of the PLAA with GRPE, Corollary 2 states that LDHD generalization can be achieved if and only if the target concept can be represented by the elements in ${\mathcal{B}}_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}$ that have a dependence on ${\mathcal{X}}_{N_{0}}$.
@@
Consider the concept $c(x)=f_{\text{PLAA}}^{\text{GRPE},{\mathcal{U}}}(x;p^{*})$:
$\sum_{k=1}^{r}c_{k}\sum_{1\leq i\leq j\leq N}(U_{k})_{ij}b_{ij}^{\text{PLAA}}(x).$
Under the conditions of Theorem 4, the PLAA with GRPE achieves LDHD generalization from ${\mathcal{X}}_{N_{0}}$ to ${\mathcal{X}}_{N}$ if and only if
@@
While the projection in the RFM and the position embeddings in the PLAA may introduce an inductive bias that benefits LDHD generalization, they can reduce the expressiveness of the models. The point is to align the models with the concept class, incorporating a strong inductive bias while maintaining sufficient expressiveness.
Figure 2: Length generalization of Transformer with RPE and RPE-Square in the unaligned copy and the URF addition tasks. Unaligned Copy: Transformers with RPE (a) and RPE-Square (b) are trained on lengths 1–5 for 1000 steps and tested on lengths 1–10. While both models generalize in-distribution, only RPE-Square achieves out-of-distribution generalization for lengths 6–10. URF Addition: Models are trained on URF 4-addition and tested on URF 5-addition. (c) Both the models are trained for $20000$ steps. The comparison result shows that the RPE fails while the RPE-Square succeeds in achieving length generalization. (d) RPE-Square trained for 200,000 steps achieves nearly perfect accuracy, with digitwise accuracy shown for each $z_{k}$.
5 Further Implications
We discuss further implications of the LDHD generalization perspective for practical length generalization techniques.